Kara spent 16.00 on tubes of pain and brushes for art class. Each tube of paint cost $3.00 and each brush costs .50. Kara purchases twice as many brushes as tubes of paint. Find the number of burshes and the number of tubes of pain that she purchased.
Answer by ankor@dixie-net.com(22740)

Kara spent 16.00 on tubes of pain and brushes for art class.
:
I didn't know you had to pay for pain, thought it just came on it's own!
:
Anyway
Each tube of paint cost $3.00 and each brush costs .50.
3p + .50b = 16
:
Kara purchases twice as many brushes as tubes of paint.
b = 2p
:
Find the number of brushes and the number of tubes of pain that she purchased.
:
Replace b with 2p in the 1st equation
3p + .50(2p) = 16
3p + 1P = 16
4p = 16
p = 16/4
p = 4 tubes of pain
then
b = 2(4)
b = 8 brushes
:
:
Check this: 3(4) + .5(8) = 16